OpenStudy (anonymous):

WILL GIVE MEDAL IF YOU HELP ME! y + 6 = -4(x - 3) B. y + 3 = -4(x - 6) C. y - 2 = -4(x - 1) D. y - 1 = -4(x - 2) LINE IS (1,2) AND (3,-6) AND SLOPE IS -4

OpenStudy (anonymous):

well just plug in the x and the y into the equation

OpenStudy (anonymous):

x is 1 and y 2 for the other one x is 3 y is -6

OpenStudy (anonymous):

your showing the same questions over and over @xxxghostxxx please try one on your own it is not that hard.

OpenStudy (anonymous):

so just plug it in the equation

OpenStudy (anonymous):

and see which one is matching with yours

OpenStudy (anonymous):

d and b?

OpenStudy (anonymous):

how you get d and b

OpenStudy (anonymous):

you put it in the wrong way

OpenStudy (anonymous):

so do you get it now

OpenStudy (anonymous):

a and c

OpenStudy (anonymous):

y - y1 = m(x - x1) plug it into the x1s and y1s

OpenStudy (anonymous):

x1 = x and y1 = y so the point (1,2) AND (3,-6) x = 1 and 3 and y = 2 and -6

OpenStudy (anonymous):

so what will the answers be?

OpenStudy (anonymous):

y-2=-4(x-1) and y-6=-4(x+3)

OpenStudy (anonymous):

so it is the 1st one and the 3rd one

OpenStudy (anonymous):

thks man

OpenStudy (anonymous):

yw